# Identifying biological perturbation targets through causal differential networks

### v2

@@ B.1 Cdn training
We trained until convergence (no improvement for 50 epochs on synthetic; 10 epochs on Perturb-seq) and selected the best model based on validation mAP (Figure 4). The validation distribution was identical to training and did not contain test causal mechanisms. During training, we used 15 CPU workers (primarily for local graph estimates) and 1 A6000 GPU. Cdn took around 6 hours to train on synthetic data, and 1 hour to finetune on real data. Unless otherwise noted, the causal structure learner was initialized from the Sea inverse covariance, FCI pretrained weights (Wu et al., 2024).
@@ Appendix C Theoretical motivation
In this analysis, we assume that the “causal” representations $h$ contains enough information both to retain global statistics and recover the true graph $E$. The former is reasonable due to high model capacity, and the latter is based on high empirical performance in graph reconstruction (Wu et al., 2024). We emphasize that the latter is an empirical judgment, which may not hold on all datasets in practice.
@@ Appendix E Additional analyses
Due to space limitations, we report uncertainty estimates in Table 7. Multiple baselines produce deterministic results (Linear, Dge, GenePt), so instead of model randomness, we report uncertainties over the sampling of single cells. Specifically, for each perturbation with $M$ cells, we sample
@@
cells uniformly at random, repeated 5 times.

### v4

@@ B.1 Cdn training
We trained until convergence (no improvement for 50 epochs on synthetic; 10 epochs on Perturb-seq) and selected the best model based on validation mAP (Figure 4). The validation distribution was identical to training and did not contain test causal mechanisms. During training, we used 15 CPU workers (primarily for local graph estimates) and 1 A6000 GPU. Cdn took around 6 hours to train on synthetic data, and 1 hour to finetune on real data. Unless otherwise noted, the causal structure learner was initialized from the Sea inverse covariance, FCI pretrained weights (Wu et al., 2025).
@@ Appendix C Theoretical motivation
In this analysis, we assume that the “causal” representations $h$ contains enough information both to retain global statistics and recover the true graph $E$. The former is reasonable due to high model capacity, and the latter is based on high empirical performance in graph reconstruction (Wu et al., 2025). We emphasize that the latter is an empirical judgment, which may not hold on all datasets in practice.
@@ D.1 Biological datasets
Top differentially expressed genes
Here, we limited our analysis to the top 1000 DEGs. In the literature, it is quite common to restrict analysis to subsets of genes, though the threshold and criteria may vary (Luecken & Theis, 2019). Selecting the top 1000 genes (based on differential expression p-value) is reasonably permissive, since these genes are selected per perturbation. For a sense of scale, Nadig et al. (2024) writes:
“In a genome-scale Perturb-seq screen, we find that a typical gene perturbation affects an estimated 45 genes, whereas a typical essential gene perturbation affects over 500 genes.”
A classic approach for identifying drug targets argues that targets are not highly differentially expressed, but network-based approaches are useful for predicting them from gene expression (Isik et al., 2015).This work also samples 1000 candidate targets as decoys.
Gene co-expression
@@ Appendix E Additional analyses
Due to space limitations, we report uncertainty estimates in Tables 7 and 8. Multiple baselines produce deterministic results (Linear, Dge, GenePt), so instead of model randomness, we report uncertainties over the sampling of single cells (Table 7). Specifically, for each perturbation with $M$ cells, we sample
@@
cells uniformly at random, repeated 5 times. We also consider the robustness of our model to samplings of the test set (Table 8). Typically, the largest cluster (assigned to train, e.g. nearly 1/3 of K562) contains weak perturbations with minimal phenotype, so different clusterings don’t introduce significant heterogeneity. Instead, we can partition all perturbations uniformly into 5 sets for the unseen cell line setting, across which the variation in results is minimal.
@@
Table 11 reports results on larger graphs. There is some drop in performance, especially Polynomial. This could be attributed to error propagation from the causal graph predictor, which performs less well (at predicting graphs) on larger graphs, and could potentially be ameliorated by finetuning on larger synthetic data (similar to how Perturb-seq models were finetuned to be larger). Linear + Soft remains quite good, perhaps because the theory (Appendix C) suggests that the summary statistics are sufficient for predicting targets, without the graph prediction.
@@ 20 Ut-Igsp 1 78 19 26
Table 8: To quantify the uncertainty across draws of the test set (perturbations, instead of cells), we partition all perturbations uniformly into 5 sets for the unseen cell line setting on K562 genome-wide. The variation across test sets is minimal.
Test Split rank $\rho$ $r$ Recall@1 Recall@5 Recall@20
Original ($n=678$) 0.974 0.915 0.921 0.699 0.900 0.957
New splits ($n=353,\times 5$) 0.973 ±$\scriptstyle\pm 0.009$ 0.917 ±$\scriptstyle\pm 0.017$ 0.923 ±$\scriptstyle\pm 0.016$ 0.696 ±$\scriptstyle\pm 0.031$ 0.908 ±$\scriptstyle\pm 0.032$ 0.961 ±$\scriptstyle\pm 0.021$
@@
Table 11: While the majority of classic causal discovery models struggle to scale to larger graphs, we show that Cdn still performs reasonably on $N=100,E=200$ graphs. AUC reported over 10 i.i.d. graphs.
Intervention Linear (1) Linear (3) Sigmoid (1) Sigmoid (3) Polynomial (1) Polynomial (3)
Hard .$64\pm.04$$65\pm.08$$63\pm.02$$61\pm.06$$50\pm.01$$49\pm.03$
Soft .$92\pm.01$$90\pm.03$$72\pm.02$$73\pm.02$$53\pm.03$$52\pm.05$